OpenStudy (kamille):

I need to find x: (a+x)/(b-x)=(4a)/(3b) Anyone has any ideas?

OpenStudy (kamille):

Also, I know that \[b \neq 0 \] and \[b \neq x\]

OpenStudy (kamille):

\[(\frac{ a+x }{ b-x } )=\frac{ 4a }{ 3b }\]

OpenStudy (kropot72):

First multiply both sides by (b - x) to eliminate the fraction of the left hand side. Then multiply both sides of the result by 3b to eliminate the fraction on the right hand side. Can you do that?

OpenStudy (kamille):

Well, I get this. If you have time, can you look? (a+x)/(b-x)=(4a/3b)|*(b-x) a+x=4a/3b(b-x)|*3b 3ba+3bx=4a/b-x

OpenStudy (kamille):

If you wont understand what I get, i can use "equation" to show.

OpenStudy (kropot72):

Multiplying both sides by (b - x) gives: \[\frac{a+x}{b-x}\times (b-x)=\frac{4a(b-x)}{3b}\] \[a+x=\frac{4a(b-x)}{3b}\] Now multiplying both sides by 3b gives: \[3b(a+x)=4a(b-x)\] Do you follow so far?

OpenStudy (kamille):

Well, I understand this. What to do next?

OpenStudy (kropot72):

Next multiply out to remove the brackets, then rearrange to get the two terms in x on the left hand side.

OpenStudy (kropot72):

Finally you factorise out x and perform a division to get: \[x=\frac{?}{?}\]

OpenStudy (kamille):

Oh,thank you very much. 3ba+3bx=4ab-4ax 3bx=4ab-3ab-4ax 3bx+4ax=ab x(3b+4a)=ab x=ab/3b+4a correct?

OpenStudy (kropot72):

Good work! Your answer is correct :)

OpenStudy (kamille):

Thank you very much:))

OpenStudy (kropot72):

You're welcome :)